OpenStudy (anonymous):

expand the logarithm. simplify if possible. log base 2 of x/y

hartnn (hartnn):

\(\large \log(A/B) = \log A -\log B\) so what about log (x/y) =... ?

OpenStudy (anonymous):

i have no idea how these work:/

hartnn (hartnn):

 in words, log of division = difference of logs